OpenStudy (anonymous):

find the inverse function of y = 2x^2 - 8x

OpenStudy (anonymous):

\[2x ^{2}-8x\]

OpenStudy (anonymous):

the answer (i think) should be \[(8+\sqrt{64+8x})/4\] through use of quadratic equation and solving for "root" but i want to know why...

OpenStudy (anonymous):

To find the inverse function, u must make x the subject of the equation. y = 2x^2 - 8x so, start by factorising, y= 2[x^2 -4x] then note that (x-2)^2 = x^2 -4x+4, note that it is similar to the top, just missing the variable +4 so we go ahead to add that in, y= 2[x^2 -4x +4 -4] (note, u must put in +4 and -4 to make the eqn still equal.) y= 2[ (x-2)^2 -4] now u can easily arrange it to make x the subject, (y/2) +4 = (x-2)^2 \[\pm \sqrt{y/2 +4}=x-2 \] \[x=2\pm \sqrt{y/2+4}\] Interchange both x and y, \[y=2\pm \sqrt{x/2+4}\] which is equal to the ans u have, just more simplified.

OpenStudy (anonymous):

thank you!!!

OpenStudy (anonymous):

Np!